Lilian9667:

What is the value of r? 1.9r+0.1r=0.6 Enter your answer as a decimal in the box. r =

Lilian9667:

This confusing?

xxcoffee:

yes

sylvester:

1.9r+0.1r=0.6 2r = 0.6 r = 0.6 / 2 r = 0.3

Lilian9667:

Hmm I think I kinda Get it?

dericawarren:

i jut had this question hold on

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dericawarren i jut had this question hold on \(\color{#0cbb34}{\text{End of Quote}}\) Ohh You did

Mathew222:

kk

Mathew222:

what is it

Mathew222:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 \(\color{#0cbb34}{\text{Originally Posted by}}\) @dericawarren i jut had this question hold on \(\color{#0cbb34}{\text{End of Quote}}\) Ohh You did \(\color{#0cbb34}{\text{End of Quote}}\)

dericawarren:

yes

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mathew222 what is it \(\color{#0cbb34}{\text{End of Quote}}\) *Clears throat* What ya think

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sylvester 1.9r+0.1r=0.6 2r = 0.6 r = 0.6 / 2 r = 0.3 \(\color{#0cbb34}{\text{End of Quote}}\) It's 0.3

dericawarren:

hold on

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dericawarren hold on \(\color{#0cbb34}{\text{End of Quote}}\) ok

dericawarren:

it want let me go back i am srry

Bulldogfan1234:

r is 0.3

Lilian9667:

Nvm, it is lol

Lilian9667:

Thank you'

jeweldiaz130:

6/91 or 0.06